# Neural Deconstruction Search for Vehicle Routing Problems

### v1

@@ 3.4 Search
The improvement step, the core component of the overall search algorithm, is depicted in Figure 2. The process begins with an initial solution $s_{0}$ that is passed to the policy DNN, which generates $K$ rollouts, each consisting of $M$ actions that specify the customers to be removed. Once the policy DNN completes its execution, these rollouts are sequentially applied to produce new candidate solutions. Specifically, the solution $s_{0}$ is first deconstructed based on the actions from the first rollout (yielding $\bar{s_{0}}$) and then reconstructed into $s^{\prime}_{0}$. After reconstruction, a simulated annealing (SA) based acceptance criterion is used to determine whether $s^{\prime}_{0}$ or $s_{0}$ should be retained, resulting in $s_{1}$. This process is repeated in each subsequent iteration. After $K$ iterations, the final solution $s_{K}$ is returned, representing the outcome of $K$ consecutive deconstruction and reconstruction operations. By performing these iterations sequentially, the solution $s_{0}$ is significantly modified, often leading to notable cost improvements between the initial input $s_{0}$ and the final output $s_{K}$.
@@
We introduced a novel simulated annealing (SA) algorithm to conduct a high-level search specifically designed for GPU-based parallelization. While parallel SA algorithms have been proposed in prior work, (Ferreiro et al., 2013; Jeong & Kim, 1990; Onbaşoğlu & Özdamar, 2001), their main concern is on the information exchange between CPU or GPU cores. In contrast, our approach focuses on executing parallel rollouts of the policy network on the GPU.
@@
[…]
@@
[…]
@@
[…]
@@ 4 Experiments
We evaluate NDS on three VRP variants with 100 to 2000 customers and compare to state-of-the-art learning-based and traditional OR methods. Additionally, we provide ablation experiments for the individual components of NDS and evaluate the generalization across different instance distributions. All experiments are conducted on a research cluster utilizing a single Nvidia A100 GPU per run. We will release our implementation of NDS, along with the instance generators, under an open-source license upon acceptance.
4.1 Problems
@@
The CVRP is one of the most extensively studied variants of the VRP. The goal is to determine the shortest routes for a fleet of vehicles tasked with delivering goods to a set of $N$ customers. Each vehicle begins and ends its route at a depot and is constrained by a maximum carrying capacity. We use the instance generator from Kool et al. (2019) to create scenarios with uniformly distributed customer locations, and the generator from Queiroga et al. (2022) for generating more realistic instances, with clustered customer locations to better simulate real-world conditions.
@@
The VRPTW extends the traditional CVRP by adding time constraints for customer deliveries. Each customer has a time window, defining the earliest and latest allowable delivery times. Vehicles can arrive early but must wait until the window opens, adding scheduling complexity. All routes start at a central depot, with a fixed service duration for deliveries and travel times based on the Euclidean distance. The objective is to minimize the total travel time while respecting both vehicle capacity and time windows, making VRPTW more complex than the standard CVRP. To generate customer locations and demands, we use the CVRP instance generator from Queiroga et al. (2022), while time windows are generated following the methodology outlined by Solomon (1987).
@@
The PCVRP is a variant of the VRP in which not all customers need to be visited. Each customer is associated with a prize, and the objective is to minimize the total travel cost minus the sum of collected prizes. Similar to the CVRP, all vehicles start and end their routes at a central depot and are constrained by vehicle capacities. To generate PCVRP instances, we use the instance generator from Queiroga et al. (2022) to create customer locations and demands. Customer prize values are generated at random, with higher prizes assigned to customers with greater demand, reflecting the increased resources required to service them.
@@ 4.2 Search Performance
We compare NDS to several heuristic solvers, including HGS (Vidal, 2022), SISRs (Christiaens & Vanden Berghe, 2020), and LKH3 (Helsgaun, 2017). Additionally, we include PyVRP (Wouda et al., 2024) (version 0.9.0), which is an open-source extension of HGS for other VRP variants. For the CVRP, we further compare NDS to the state-of-the-art learning-based methods, SGBS-EAS (Choo et al., 2022), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), and GLOP (Ye et al., 2024b).
@@
For each problem and problem size, we perform a separate training run. Training consists of $2000$ epochs for settings with 1000 or fewer customers. For the 2000 customer setting, we resume training from the 1000 customer model checkpoint at 1500 epochs and train for an additional $500$ epochs. In each epoch, we process 1500 instances, with each instance undergoing $100$ iterations, $128$ rollouts, and $10$ initial improvement steps. The learning rate is set to $10^{-4}$ and $15$ customers are selected per deconstruction step across all problem sizes. The training durations are approximately 5, 8, 15, and 8 days for the problem sizes 100, 500, 1000 and 2000, respectively. The training curves are presented in Appendix A, while visualizations of policy rollouts are available in Appendix B.
@@
At test time, we limit the runtime to $5$, $60$, $120$, and $240$ seconds of wall time per instance for HGS, SISRs, and NDS to ensure a fair comparison, as these methods process test instances sequentially. SGBS-EAS and LEHD, which process instances in batches, are given an equivalent search budget per batch. All approaches are restricted to using a single CPU core. For the CVRP, we use the test instances from Kool et al. (2019) for $N$=$100$ (10,000 instances), Drakulic et al. (2023) for $N$=$500$ (128 instances), and Ye et al. (2024b) for $N$=$1000$ and $N$=$2000$ (100 instances each). For the VRPTW and PCVRP, we generate new test sets consisting of 10,000 instances for $N$=$100$ and 250 instances for settings with more than 100 customers.
@@
[…]
Results
Table 1 presents the performance of all compared methods on the test data. The gap is reported relative to HGS for the CVRP, and to PyVRP-HGS for the VRPTW and PCVRP. Across the 12 test settings, NDS delivers the best performance in 11 cases, with HGS being the only approach able to outperform it on the CVRP with 100 customers. Compared to other learning-based methods, NDS shows significant performance improvements across all CVRP sizes. On the CVRP with 2000 customers, NDS achieves a 7 percentage point improvement over the best-performing learning-based method, LEHD, and a 12 percentage point improvement over GLOP. Against the state-of-the-art HGS and its extension, PyVRP-HGS, NDS performs especially well on larger instances, achieving a gap of more than 2% across all problems for instances with 2000 customers. For the PCVRP, NDS also attains substantial gaps relative to PyVRP-HGS, exceeding 4% on instances with 500 or more nodes. When compared to SISRs, NDS maintains a small advantage on larger instances and demonstrates significantly better performance on small instances.
Table 1: Performance on test data. The gap is calculated relative to HGS for the CVRP and relative to PyVRP-HGS for the VRPTW and PCVRP. Runtime is reported on a per-instance basis in seconds.
@@ Results
LKH3333 15.64 0.50% 41 37.25 1.66% 174 42.16 1.61% 408 58.12 1.35% 1448
SGBS-EAS 15.59 0.17% 5 - - - - - - - - -
BQ (BS64) 15.74 1.13% 1 37.51 2.32% 23 43.32 4.36% 164 - - -
LEHD (RRC) 15.61 0.30% 5 37.04 1.04% 60 42.47 2.31% 121 60.11 4.76% 246
GLOP (LKH3) - - - - - - 45.90 10.58% 4 63.00 9.79% 6
@@ 4.3 Ablation Studies
We perform a series of ablation experiments to assess the importance of different components of NDS. These experiments are conducted on separate validation instances with $N$=$500$ customers. The parameter configuration remains identical to the previous section, except the training is reduced to 1,000 epochs and the ASA search is limited by the number of iterations. For the CVRP and VRPTW, we run 1,000 iterations using 8 augmentations, while for the PCVRP, we perform 50 iterations with 128 augmentations.
@@
Table 2: Ablation experiments.
@@
Insertion Order
The insertion algorithm reinserts removed customers in a specified order. During testing, we reconstruct a deconstructed solution five times using different customer orders and retain the best solution. For the first reconstruction iteration, we use the customer order provided by the DNN, while for the remaining four iterations we use a random order. We compare our standard setting to using only random orderings for all five insertion iterations to assess whether the ordering enhances overall search performance. The results in Table 2b show that using a only random orderings leads to significantly worse performance across all three problems, indicating that the learned policy not only plays a crucial role in deconstruction, but also significantly influences reconstruction.
Learned Policy
We assess the relevance and effectiveness of the learned deconstruction policy by replacing it with a handcrafted heuristic based on the destroy procedure outlined in Christiaens & Vanden Berghe (2020). The resulting approach eliminates any learned components, but is otherwise identical to NDS. The performance comparison, shown in Figure 2c, reveals that the heuristic deconstruction policy performs significantly worse than the learned counterpart, with performance gaps of up to 1.5% on the PCVRP. This demonstrates that the DNN is capable of learning a highly efficient policy that surpasses handcrafted methods in this use case.
@@ 4.4 Generalization
One major advantage of learning-based solution approaches is their ability to adapt precisely to the specific type of instances at hand. However, in real-world scenarios, concept drift in the instance distributions cannot always be avoided. In this experiment, we evaluate whether the learned policies of NDS can handle instances sampled from a slightly different distribution. For the CVRP with $N$=$500$, we train a policy on instances with medium-capacity vehicles and customer locations that follow a mix of uniform and clustered distributions. We then evaluated the learned policy on instances with low- and high-capacity vehicles, and customer locations following either uniform or clustered distributions. Additionally, we train distribution-specific models for each test setting for comparison. As a baseline, we compare against HGS and SISRs, giving all approaches the same runtime. The results are shown in Table 3, where NDS (OOD) represents the model’s performance when the training and test distributions differ, and NDS (ID) represents the setting where the training and test distributions are identical. Overall, the performance difference between the two settings is minimal, indicating that NDS generalizes well across different distributions. Interestingly, the distribution of customer locations has a larger impact on performance than vehicle capacity.
@@ 5 Conclusion
In this work, we introduced a novel search method, NDS, which leverages a learned policy to deconstruct solutions for routing problems. NDS presents several key advantages. First, it delivers superior performance, consistently outperforming state-of-the-art OR methods under equal runtime. Second, NDS scales effectively to larger problem instances, handling up to $N$=$2000$ customers, due to the fact that the number of customers selected by the policy is independent of the problem size. Third, it demonstrates strong generalization across different data distributions. Finally, NDS is easily adaptable to new vehicle routing problems, requiring only small adjustments to the greedy insertion heuristic and the model input.
@@
A notable limitation is the reliance on a GPU for executing the policy network. Future research could explore model distillation techniques to lower the computational requirements or investigate whether the underlying principles of the learned policies can be approximated using faster, more efficient algorithms.
Acknowledgements
@@
André Hottung was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Grant No. 521243122. Paula Wong-Chung was supported by the German Academic Exchange Service Research Internships in Science and Engineering (DAAD RISE) program and by the Globalink Research Internship (GRI) program. Additionally, we gratefully acknowledge the funding of this project by computing time provided by the Paderborn Center for Parallel Computing (PC2). Furthermore, some computational experiments in this work have been performed using the Bielefeld GPU Cluster. We thank the HPC.NRW team for their support.
Appendix A Training Curves

### v2

@@ 4.4 Search
The improvement step, the core component of the overall search algorithm, is depicted in Figure 2. Note that the improvement step procedure is designed to limit the data transfer between GPU and CPU while exploring a large number of different solutions. The process begins with an initial solution $s_{0}$ that is passed to the policy DNN, which generates $K$ rollouts, each consisting of $M$ actions that specify the customers to be removed. Once the policy DNN completes its execution, these rollouts are sequentially applied to produce new candidate solutions. Specifically, the solution $s_{0}$ is first deconstructed based on the actions from the first rollout (yielding $\bar{s_{0}}$) and then reconstructed into $s^{\prime}_{0}$. After reconstruction, a simulated annealing (SA) based acceptance criterion is used to determine whether $s^{\prime}_{0}$ or $s_{0}$ should be retained, resulting in $s_{1}$. This process is repeated in each subsequent iteration. After $K$ iterations, the final solution $s_{K}$ is returned, representing the outcome of $K$ consecutive deconstruction and reconstruction operations. By performing these iterations sequentially, the solution $s_{0}$ is significantly modified, often leading to notable cost improvements between the initial input $s_{0}$ and the final output $s_{K}$, while only needing to transfer data from the GPU to the CPU a single time.
@@
We introduced a novel simulated annealing (SA) algorithm to conduct a high-level search specifically designed for GPU-based parallelization. While parallel SA algorithms have been proposed in prior work, (Ferreiro et al., 2013; Jeong & Kim, 1990; Onbaşoğlu & Özdamar, 2001), their main concern is on the information exchange between CPU cores. In contrast, our approach focuses on executing parallel rollouts of the policy network on the GPU.
@@
1:procedure Search(Instance $l$, Number of iterations $\mathit{maxIter}$, number of augmentations $A$, number of rollouts $K$, start temperature $\lambda^{start}$, temperature decay rate $\lambda^{decay}$, trained policy network $\pi_{\theta}$, threshold factor $\delta$)
@@
7:    $\mathit{cost_{a}}\leftarrow\textsc{Obj}(s_{a})\qquad\forall a\in\{1,\dots,A\}$
8:    $\mathit{cost^{*}}\leftarrow\min(\mathit{cost_{0}},\dots,\mathit{cost_{A}})$
9:    $\mathit{thresh}\leftarrow\mathit{cost^{*}}+(\lambda\times\delta)$
@@
11:     if $\mathit{cost_{a}}>\mathit{thresh}$ then
12:      $s_{a}\leftarrow\textsc{RandomChoice}(\{s^{\prime}\in\{s_{0},\dots s_{A}\}\mid\textsc{Obj}(s^{\prime})<\mathit{thresh}\})$
@@ 5 Experiments
We evaluate NDS on three VRP variants with 100 to 2000 customers and compare to state-of-the-art learning-based and traditional OR methods. Additionally, we provide ablation experiments for the individual components of NDS and evaluate the generalization across different instance distributions. All experiments are conducted on a research cluster utilizing a single Nvidia A100 GPU per run. Our implementation of NDS is available at https://github.com/ahottung/NDS.
5.1 Problem Instances
@@
We use the instance generator from Kool et al. (2019) to create scenarios with uniformly distributed customer locations, and the generator from Queiroga et al. (2022) for generating more realistic instances with clustered customer locations to better simulate real-world conditions.
@@
We use the instance generator from Queiroga et al. (2022) to generate customer locations and demands, while time windows are generated following the methodology outlined by Solomon (1987).
@@
To generate PCVRP instances, we use the instance generator from Queiroga et al. (2022) to create customer locations and demands. Customer prize values are generated at random, with higher prizes assigned to customers with greater demand, reflecting the increased resources required to service them.
@@ 5.2 Search Performance
We compare NDS to several heuristic solvers, including HGS (Vidal, 2022), SISRs (Christiaens & Vanden Berghe, 2020), and LKH3 (Helsgaun, 2017). Additionally, we include PyVRP (Wouda et al., 2024) (version 0.9.0), which is an open-source extension of HGS for other VRP variants. For the CVRP, we further compare NDS to the state-of-the-art learning-based methods, SGBS-EAS (Choo et al., 2022), NeuOpt (Ma et al., 2023), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), UDC (Zheng et al., 2024), and GLOP (Ye et al., 2024b). We run all baselines ourselves on the same test instances. More details on the baseline settings can be found in Appendix A.
@@
For each problem and problem size, we perform a separate training run. Training consists of $2000$ epochs for settings with 1000 or fewer customers. For the 2000 customer setting, we resume training from the 1000 customer model checkpoint at 1500 epochs and train for an additional $500$ epochs. In each epoch, we process 1500 instances, with each instance undergoing $100$ iterations, $128$ rollouts, and $10$ initial improvement steps. The learning rate is set to $10^{-4}$ and $15$ customers are selected per deconstruction step across all problem sizes. The training durations are approximately 5, 8, 15, and 8 days for the problem sizes 100, 500, 1000 and 2000, respectively. The training curves are presented in Appendix B, while visualizations of policy rollouts are available in Appendix C.
@@
At test time, we limit the runtime to $5$, $60$, $120$, and $240$ seconds of wall time per instance for HGS, SISRs, and NDS to ensure a fair comparison, as these methods process test instances sequentially. For most learning-based approaches we report results for two different termination criteria, i.e., the original termination criteria proposed in the respective paper and a new wall-time-based termination criteria. For approaches that process instances in batches, an adjusted time limit is applied to each batch as a whole, ensuring that the total time required to solve the complete set matches that of approaches using a per-instance time limit. Note that this experimental setup favors batch-based approaches over NDS, as processing instances in batches is computationally more efficient, but not feasible in all applications. All approaches are restricted to using a single CPU core. For the CVRP, we use the test instances from Kool et al. (2019) for NN=$100$ (10,000 instances), Drakulic et al. (2023) for NN=$500$ (128 instances), and Ye et al. (2024b) for NN=$1000$ and NN=$2000$ (100 instances each). For the VRPTW and PCVRP, we generate new test sets consisting of 10,000 instances for NN=$100$ and 250 instances for settings with more than 100 customers. To allow for a fair comparison, we ensure that all reinforcement learning based approaches are trained on instances from the same distribution that is also used for sampling the test instances.
@@
For NDS, the starting temperature $\lambda^{\mathit{start}}$ is set to $0.1$ and decays exponentially to $0.001$ throughout the search. The threshold factor $\delta$ is fixed at $15$. During the improvement step, $200$ rollouts are performed per instance, and each deconstructed solution is reconstructed $5$ times ($1\times$ based on the selected order of the DNN and $4\times$ using a random customer order). The number of augmentations is set to $8$ for the CVRP and VRPTW, and $128$ for the PCVRP. Note that this configuration was manually selected rather than derived from automated hyperparameter tuning. In Appendix E, we analyze the impact of the hyperparameters on search performance.
Table 1: Performance on test data. The gap is calculated relative to HGS for the CVRP and relative to PyVRP-HGS for the VRPTW and PCVRP. Runtime is reported on a per-instance basis in seconds. The best results (i.e., those with the lowest objective function value) are shown in bold, and the second-best are underlined.
@@
LKH33 15.64 0.50% 41 37.25 1.66% 174 42.16 1.61% 408 58.12 1.35% 1448
SGBS-EAS   Iter. Limit 15.59 0.20% 2 - - - - - - - - -
                  Time Limit 15.59 0.17% 5 - - - - - - - - -
NeuOpt          10k Iter. 15.66 0.59% 1 - - - - - - - - -
                  Time Limit 15.59 0.15% 5 - - - - - - - - -
BQ        Beam width 16 15.81 1.55% 1 37.64 2.68% 10 43.53 4.86% 45 61.75 7.61% 323
          Beam width 64 15.74 1.13% 1 37.51 2.32% 23 43.32 4.36% 164 - - -
LEHD           1000 Iter. 15.63 0.41% 1 37.10 1.21% 28 42.41 2.17% 159 59.45 3.60% 1476
                  Time Limit 15.61 0.30% 5 37.04 1.04% 60 42.47 2.31% 121 60.11 4.76% 246
UDC           250 Stages - - - 37.71 2.86% 13 42.77 3.04% 21 - - -
                  Time Limit - - - 37.63 2.67% 60 42.65 2.76% 121 - - -
GLOP (LKH3) - - - - - - 45.90 10.58% 1 63.02 9.82% 2
@@
Results
Table 1 presents the performance of all compared methods on the test data. The gap is reported relative to HGS for the CVRP, and to PyVRP-HGS for the VRPTW and PCVRP. By choosing HGS for the gap calculation, we follow recent publications and ensure easy comparability of our results with these earlier works. Across the 12 test settings, NDS delivers the best performance in 11 cases, with HGS being the only approach able to outperform it on the CVRP with 100 customers. Compared to other learning-based methods, NDS shows significant performance improvements across all CVRP sizes. On the CVRP with 2000 customers, NDS achieves a 6.83% improvement over the best-performing learning-based method, LEHD (when both are given the same runtime), and an 11.13% improvement over GLOP. Against the state-of-the-art HGS and its extension, PyVRP-HGS, NDS performs especially well on larger instances, achieving an improvement of more than 2% across all problems for instances with 2000 customers. For the PCVRP, NDS also attains substantial improvements relative to PyVRP-HGS, exceeding 4% on instances with 500 or more nodes. When compared to SISRs, NDS maintains a small advantage on larger instances and demonstrates significantly better performance on small instances.
5.3 Performance Across Different Runtime Limits
To better understand the application scenarios of NDS, we evaluate its performance under varying runtime limits. In this experiment, we compare NDS only to SISRs as it is the overall best performing baseline in the previous experiment (see Section 5.2). We use the same trained models and evaluation setup as in the previous experiment, but with test sets reduced to 100 instances each to manage computational costs.
Figure 3 presents the results for instances with 500 customers. For each problem, five different runtime limits ranging from 15 to 250 seconds are considered. The hyperparameters of SISRs and NDS remain consistent with those used in previous experiments across all runtime limits. Notably, NDS outperforms SISRs across all runtime limits, except for the PCVRP, where SISRs perform better under a 15-second runtime limit. The relative performance gap is particularly striking for the CVRP, where running NDS for just 30 seconds surpasses the performance of SISRs at 250 seconds. Similar trends are observed for instances with 100, 1000, and 2000 customers. Additional details and results for other problem sizes are provided in Appendix D.
Figure 3: Comparison of NDS and SISRs performance across varying runtime limits.
@@ 5.4 Ablation Studies
We perform a series of ablation experiments to assess the importance of different components of NDS. These experiments are conducted on separate validation instances with NN=$500$ customers. The parameter configuration remains identical to the previous section, except the training is reduced to 1,000 epochs and the ASA search is limited by the number of iterations. For the CVRP and VRPTW, we run 1,000 iterations using 8 augmentations, while for the PCVRP, we perform 50 iterations with 128 augmentations. In addition to the ablation results reported here, we present an ablation study on the impact of the initial improvement step during training in Appendix F.
@@
Insertion Order
The insertion algorithm reinserts removed customers in a specified order. During testing, we reconstruct a deconstructed solution five times using different customer orders and retain the best solution. For the first reconstruction iteration, we use the customer order provided by the DNN, while for the remaining four iterations we use a random order. We compare this standard setting to using only random orders across all five insertion iterations to assess whether the DNN has learned to select an order that improves the overall search performance. The results in Table 2b show that using a only random orderings leads to significantly worse performance across all three problems, indicating that the learned policy not only plays a crucial role in deconstruction, but also significantly influences reconstruction.
Learned Policy
We assess the relevance and effectiveness of the learned deconstruction policy by replacing it with a handcrafted heuristic based on the destroy procedure outlined in Christiaens & Vanden Berghe (2020). The resulting approach eliminates any learned components, but is otherwise identical to NDS. The performance comparison, shown in Figure 2c, reveals that the heuristic deconstruction policy performs significantly worse than the learned counterpart, with performance gaps of up to 1.5% on the PCVRP. This demonstrates that the DNN is capable of learning a highly efficient policy that surpasses handcrafted methods.
Table 2: Ablation experiments. All tables report objective function values.
@@ 5.5 Generalization
One major advantage of learning-based solution approaches is their ability to adapt precisely to the specific type of instances at hand. However, in real-world scenarios, concept drift in the instance distributions cannot always be avoided. In this experiment, we evaluate whether the learned policies of NDS can handle instances sampled from a slightly different distribution. For the CVRP with NN=$500$, we train a policy on instances with medium-capacity vehicles and customer locations that follow a mix of uniform and clustered distributions. We then evaluated the learned policy on instances with low- and high-capacity vehicles, and customer locations following either uniform or clustered distributions. Additionally, we train distribution-specific models for each test setting for comparison. As a baseline, we compare against HGS and SISRs, giving all approaches the same runtime. The results are shown in Table 3, where NDS (OOD) represents the model’s performance when the training and test distributions differ, and NDS (ID) represents the setting where the training and test distributions are identical. Overall, the performance difference between the two settings is minimal, indicating that NDS generalizes well across different distributions. Interestingly, the distribution of customer locations has a larger impact on performance than vehicle capacity. In Appendix G, we also evaluate NDS’s ability to generalize to both smaller and significantly larger instances.
@@ 6 Conclusion
In this work, we introduced a novel search method, NDS, which leverages a learned policy to deconstruct solutions for routing problems. NDS presents several key advantages. First, it delivers excellent performance, consistently matching or outperforming state-of-the-art OR methods under equal runtime. Second, NDS scales effectively to larger problem instances, handling up to NN=$2000$ customers, due to the fact that the number of customers selected by the policy is independent of the problem size. Third, it demonstrates strong generalization across different data distributions. Finally, NDS is easily adaptable to new vehicle routing problems, requiring only small adjustments to the greedy insertion heuristic and the model input.
@@
A notable limitation is the reliance on a GPU for executing the policy network. Future research could explore model distillation techniques to lower the computational requirements or investigate whether the underlying principles of the learned policies can be approximated using faster, more efficient algorithms. Another limitation of NDS is its reliance on training data that closely resembles the instances encountered during testing. Future work could investigate whether fine-tuning techniques enable fast adaptation of a pretrained foundation model to new problem instances. This could significantly reduce training time and minimize the amount of required training data.
Acknowledgments
@@
André Hottung was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Grant No. 521243122. Paula Wong-Chung received funding from the Deutscher Akademischer Austauschdienst (DAAD) RISE Germany program for this research. Additionally, we gratefully acknowledge the funding of this project by computing time provided by the Paderborn Center for Parallel Computing (PC2). Furthermore, some computational experiments in this work have been performed using the Bielefeld GPU Cluster. We thank the HPC.NRW team for their support.
Appendix A Baseline Configurations
HGS
We conduct evaluation runs using the publicly available [^1] https://github.com/vidalt/HGS-CVRP implementation from the authors that has been designed and calibrated for medium-scale CVRP instances with up to 1,000 customers. We use a wall-time-based stopping criteria to limit the duration of each search run.
SISRs
Since the code of SISRs is not publicly available, we reimplement SISRs in C++ based on the description of the authors. Our implementation matches the performance reported in the original paper. We use a time based stopping criteria to limit the duration of each search run.
SGBS-EAS
We conduct evaluation runs using the code that was made publicly available [^2] https://github.com/yd-kwon/SGBS by the authors. For the CVRP with 100 customers, we use the publicly available model which has been trained on instances from the same distribution as our test instances. We run SGBS-EAS using the configuration proposed in the original paper in which search runs are limited based on the number of iterations. Additionally, we conduct experiments in which we limit the search to the same runtime as NDS. More precisely, we limit the total runtime to 13.9 hours, which corresponds to 5 seconds per instance. We increase the beam width to 7 in this experiment to increase exploration and to better use the available runtime. For the VRPTW with 100 customers, no pretrained models are publicly available. We hence train the models ourselves on instances sampled from the same distribution as our test instances to allow for fair comparison. We use the same training configuration as in the original paper (Kwon et al., 2020) and train the model for 10,000 epochs with each epoch comprising 10,000 instances. At test time, we use the same setup for the VRPTW as for the CVRP.
NeuOpt
We evaluate NeuOpt on the CVRP with 100 customers using the code and trained model made available [^3] https://github.com/yining043/NeuOpt by the authors. Note that the model has been trained on CVRP instances sampled from the same distribution as the test instances. Like SGBS-EAS, NeuOpt is evaluated only on instances with 100 customers, as its autoregressive nature makes training on larger instances impractical. We conduct two sets of experiments: one where the search is limited to 10,000 iterations (as in the original paper) and another where the search is constrained by runtime, using the same runtime limits applied to NDS. For the former, we set the dynamic data augmentation (D2A) parameter to 1, and for the latter, we set D2A to 5 to enhance exploration during the search.
BQ
We conduct evaluation runs using the code and trained model made publicly available [^4] https://github.com/naver/bq-nco by the authors. BQ is a supervised learning approach that focuses on generalization and requires near-optimal solutions during training, which are only obtainable for smaller instances. Consequently, we use the publicly available model, trained on instances with 100 customers, to solve all four problem sizes considered in our experiments. Note that the model has been trained on the same distribution as our test instances with 100 customers. We perform two sets of experiments: one with a beam width of 16 and another with a beam width of 64. In both sets, we use a batch size of 1, as the beam search constructs multiple solutions in parallel, fully utilizing the parallel computing capabilities of our GPUs. For the CVRP with 2000 customers, results are only reported for a beam width of 16, as a beam width of 64 leads to out-of-memory issues.
LEHD
We evaluate LEHD using the publicly available code [^5] https://github.com/CIAM-Group/NCO_code/tree/main/single_objective/LEHD provided by the authors. LEHD is a supervised learning approach focused on generalization and requires near-optimal solutions during training, which are only obtainable for smaller instances. Consequently, we use the publicly available model, trained on instances with 100 customers, to solve all four problem sizes considered in our experiments. Note that the model was trained on the same distribution as our test instances with 100 customers. We evaluate LEHD using the random re-construct method proposed by the authors. We conduct two sets of experiments: one where the search is limited to 1,000 iterations (as in the original paper) and another where the search is constrained by runtime, using the same runtime limits applied to NDS.
UDC
We evaluate UDC using the code that was made publicly available [^6] https://github.com/CIAM-Group/NCO_code/tree/main/single_objective/UDC-Large-scale-CO-master by the authors. For a fair comparison, we use custom models that are trained specifically for each considered problem size, instead of using the provided one-size-fits-all trained model. We use the default training configuration, but change sample_size to 30 for the CVRP with 1,000 customers to reduce GPU memory usage. For the CVRP with 2,000 customers, we are not able to conduct training due to memory constraints. The training instances are sampled from the same distribution as our test instances.
We evaluate UDC’s test-time performance in two sets of experiments: one with the search limited to 250 stages (as in the original paper) and another constrained by runtime, using the same limits applied to NDS. For the CVRP with 500 customers, a batch size of 6 is used, while for the CVRP with 1,000 customers, a batch size of 4 is applied. The batch size is chosen to fully leverage the available GPU memory.
GLOP
We evaluate GLOP using the code and trained models provided by the authors [^7] https://github.com/henry-yeh/GLOP. GLOP is a divide-and-conquer approach tailored for large-scale problems, and we evaluate it only on CVRP instances with 1,000 and 2,000 customers, as these are the instance sizes for which trained models are available. The training instances used by the original authors are sampled from the same distribution as our test instances. We run GLOP with the configuration outlined in the original paper, using LKH3 as a subsolver and a single CPU core. Notably, there is no straightforward way to extend GLOP to longer runtime settings for the CVRP.
Appendix B Training Curves